A flywheel rotates at an angular velocity (omega) = 31.4 rad / s. The moment of inertia is 63.6 kg / m². Inhibitory forces act on it. Find the braking moment M, under the action of which the flywheel stops after 20 S.


1
Expert's answer
2021-04-05T10:37:46-0400

"\tau=I\epsilon\to\tau=I\frac{\Delta\omega}{\Delta t}=63.6\cdot\frac{0-31.4}{20}=-99.85\ (N\cdot m)" . Answer